OpenStudy (anonymous):

MEDAL!! (suppose there is a group of 7 people from which we will make a committee) In how many ways can we pick a three-person committee?

jimthompson5910 (jim_thompson5910):

Let's say we have 3 slots A,B,C how many do we have to choose from for slot A?

OpenStudy (anonymous):

7?

OpenStudy (mathmate):

Correct, then for slot B?

OpenStudy (mathmate):

Recall that one of the sever is already in slot A.

OpenStudy (anonymous):

6?

jimthompson5910 (jim_thompson5910):

yep, and then C?

OpenStudy (anonymous):

5

OpenStudy (anonymous):

multiply?

jimthompson5910 (jim_thompson5910):

yes, you'll multiply 7, 6 and 5

OpenStudy (anonymous):

210

jimthompson5910 (jim_thompson5910):

that won't be your final answer but it gets you closer

jimthompson5910 (jim_thompson5910):

210 is the number of ways to pick 3 people IF order mattered but there is no ranking on this committee, so order doesn't matter

jimthompson5910 (jim_thompson5910):

so what you have to do is divide by 3! = 3*2*1 = 6 to get the correct count

jimthompson5910 (jim_thompson5910):

the reason why is because there are 6 ways to order any three objects xyz xzy yxz yzx zxy zyx

OpenStudy (anonymous):

210 divided by 3?

jimthompson5910 (jim_thompson5910):

210/6 actually

jimthompson5910 (jim_thompson5910):

or 210/(3!)

OpenStudy (anonymous):

ah okay. I read your response wrong haha so 35

jimthompson5910 (jim_thompson5910):

correct

jimthompson5910 (jim_thompson5910):

another way is to use the combination formula \[\Large _n C _r = \frac{n!}{r!*(n-r)!}\] with n = 7 and r = 3 and you'll get the same answer

OpenStudy (anonymous):

@jim_thompson5910 what if it was a four-person committee instead? wouldn't it be 7C4=35? also?

jimthompson5910 (jim_thompson5910):

that is correct

OpenStudy (anonymous):

thank you

jimthompson5910 (jim_thompson5910):

if you think of it in terms of 7 slots A through G we have this 

jimthompson5910 (jim_thompson5910):

now imagine cutting a line through that group such that one side (say the left side) has 3 slots and the other side has 4 slots 

jimthompson5910 (jim_thompson5910):

arranging 7 people to go in slots A through C is the exact same as arranging the remaining 4 to go in D through G so that's why 7 C 3 = 7 C 4 in general \[\LARGE _n C _x = \ _n C _y\] where x+y = n (so in this case, x+y = 3+4 = 7 which is the value of n)

OpenStudy (anonymous):

Yes, that is the way my instructor explains it also, thanks again!

jimthompson5910 (jim_thompson5910):

you're welcome